# Reliable algorithm selection for machine learning-guided design

### v1

@@ Appendix E RNA binder experiment details
To facilitate interpretability of the label values, following Sinai et al. [40] we normalized the ViennaFold binding energy [28] by that of the complement of the RNA target sequence, which can be seen as an estimate of the energy of the true optimal binding sequence. Consequently, a label value of $1$ means a binding energy equal to that of the complement sequence.
@@ E.1 Menu of design algorithm configurations
Predictive models.
The three predictive models were a ridge regression model, where the ridge regularization hyperparameter was set by leave-one-out cross-validation; an ensemble of three feedforward neural networks, each with two $100$-unit hidden layers; and an ensemble of three convolutional neural networks, each with three convolutional layers with $32$ filters, followed by two $100$-unit hidden layers. Each model in both ensembles was trained for five epochs using Adam with a learning rate of $10^{-3}$.
Design algorithm hyperparameter settings.
For AdaLead [40], the values of the threshold hyperparameter on the menu were $\kappa\in\{0.2,0.15,0.1,0.05,0.01\}$. For Biswas , the algorithm used by Biswas et al. [7], the values of the temperature hyperparameter on the menu were $T\in\{0.02,0.015,0.01,0.005\}$. For Conditioning by Adaptive Sampling [8], or CbAS , the values of the quantile hyperparameter on the menu were $Q\in\{0.1,0.2,\ldots,0.9\}$. For Design by Adaptive Sampling [9] ( DbAS ) with either the FF or CNN models, the values of the quantile hyperparameter on the menu were $Q\in\{0.1,0.2,\ldots,0.9\}$, and with ridge regression, $Q\in\{0.1,0.2\}$. Both CbAS and DbAS involve a generative model, which was a variational autoencoder (VAE) with $10$ latent dimensions, and feedforward models with $20$-unit hidden layers for both the encoder and decoder. Both algorithms were run for twenty iterations; each iteration retrained the VAE for five epochs using Adam with learning rate $10^{-3}$. PEX [31] was run with default values of all hyperparameters. The resulting menu of design algorithm configurations varied greatly in their the mean design prediction, as well as their deviation from the training sequences (Fig. 7).
@@ E.2 Density ratio estimation
To estimate the density ratio function, $p_{X;\lambda}(\cdot)/p_{\text{lab}}(\cdot)$, for every configuration on the menu, $\lambda\in\Lambda$, we used multinomial logistic regression-based density ratio estimation (MDRE) [44]. MDRE builds upon a formal connection between density ratio estimation (DRE) and classification [6, 22]: for the (correctly specified) binary classifier that minimizes the population cross-entropy risk in distinguishing between samples from two distributions, its logit for any input is equivalent to the density ratio at that input [22]. In practice, however, given finite samples from the two distributions, we can only hope to find a classifier that minimizes the empirical risk. This distinction between the population and empirical risks hinders DRE far more than it does classification for classification’s sake: obtaining the density ratio requires getting the exact value of the population-optimal classifier’s logit, whereas optimal classification performance can be achieved by any classifier that learns the same decision boundary as the optimal classifier, even if its logits differ for some inputs. Accordingly, we found that for many configurations on the menu, classifiers with very low training and validation losses often yielded poor approximations of the density ratio. This was particularly true when the design and labeled distributions were far apart, because the classification problem is too easy given finite samples: many different classifiers can minimize the empirical risk, none of which may happen to coincide with the population-optimal one whose logits are equivalent to density ratios. Telescoping density ratio estimation [32] tackles this problem by constructing intermediate distributions that interpolate between the numerator and denominator distributions, creating a sequence of “harder” binary classification problems for which there are fewer empirically optimal classifiers, and whose resulting estimated density ratios can be combined with a telescoping sum to approximate the original density ratio of interest. MDRE is similarly motivated but constructs a single multi-class classification problem between the intermediate distributions, justified by theoretical connections between the population-optimal classifier and the density ratio analogous to those described above [44]. Concretely, let $h^{c}(x)$ denote the unnormalized log-probability according to a trained classifier that $x$ belongs to distribution $c\in\{1,\ldots,C\}$, where $C$ denotes the total number of distributions. MDRE uses $\exp(h^{i}(x)-h^{j}(x))$ to approximate the density ratio between distributions $i$ (numerator) and $j$ (denominator).

### v2

@@ Appendix E RNA Binder Experiment Details
To facilitate interpretability of the label values, following Sinai et al., (2020) we normalized the ViennaFold binding energy (Lorenz et al.,, 2011) by that of the complement of the RNA target sequence, which can be seen as an estimate of the energy of the true optimal binding sequence. Consequently, a label value of $1$ means a binding energy equal to that of the complement sequence.
@@ E.1 Menu of Design Algorithm Configurations
See Fig. 8 for a diagram of the menu structure.
Figure 8: Menu structure for designing RNA binders. The menu for the results in the main text (size $78$) had a nested configuration space involving categorical options for the predictive model and design algorithm, and grids of real values for one hyperparameter for each design algorithm. The expanded menu for the results in Figure 10 (size 249) adds an additional real-valued hyperparameter for AdaLead , and integer-valued options for the generative model architecture for CbAS and DbAS (dash-outlined cells).
@@
Predictive models
The three predictive models were a ridge regression model, where the ridge regularization hyperparameter was set by leave-one-out cross-validation; an ensemble of three fully connected neural networks, each with two $100$-unit hidden layers; and an ensemble of three convolutional neural networks, each with three convolutional layers with $32$ filters, followed by two $100$-unit hidden layers. Each model in both ensembles was trained for five epochs using Adam with a learning rate of $10^{-3}$.
Design algorithm hyperparameter settings
For AdaLead (Sinai et al.,, 2020), the values of the threshold hyperparameter on the menu were $\kappa\in\{0.2,0.15,0.1,0.05,0.01\}$. For Biswas , the algorithm used by Biswas et al., (2021), the values of the temperature hyperparameter on the menu were $T\in\{0.02,0.015,0.01,0.005\}$. For Conditioning by Adaptive Sampling (Brookes et al.,, 2019), or CbAS , the values of the quantile hyperparameter on the menu were $Q\in\{0.1,0.2,\ldots,0.9\}$. For Design by Adaptive Sampling (Brookes and Listgarten,, 2018) ( DbAS ) with either the fully connected or convolutional models, the values of the quantile hyperparameter on the menu were $Q\in\{0.1,0.2,\ldots,0.9\}$, and with ridge regression, $Q\in\{0.1,0.2\}$. Both CbAS and DbAS involve a generative model, which was a variational autoencoder (VAE) with $10$ latent dimensions, and fully connected models with $20$-unit hidden layers for both the encoder and decoder. Both algorithms were run for twenty iterations; each iteration retrained the VAE for five epochs using Adam with learning rate $10^{-3}$. PEX (Ren et al.,, 2022) was run with default values of all hyperparameters. The resulting menu of design algorithm configurations varied greatly in their mean design prediction, as well as their distance from the training sequences (Fig. 9).
To assess how the multiplicity correction might impact selection rates for larger menus, we also ran our method with an expanded menu of $249$ configurations. In addition to the hyperparameter settings listed above, this menu included different values of the AdaLead recombination rate hyperparameter, $r\in\{0.1,0.2,0.5\}$, and numbers of hidden units ($\{5,10\}$) and latent dimensions ($\{20,100\}$) in the VAE used by both CbAS and DbAS (Fig. 8). Interestingly, our method exhibited similar error rates but higher selection rates with this larger menu, including non-zero selection rates for a broader range of $\tau$ than with the original menu of size $78$ (Fig. 10a), due to the expanded menu containing more configurations that produce higher mean design labels.
Figure 10: Design algorithm selection for designing RNA binders (expanded menu). (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for our original results (with menu of size 78; green) and our method with the expanded menu of size 249 diagrammed in Figure 8 (purple). (b) For both menus, the median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label achieved by selected configurations, over trials for which the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which the method ceased to select any configuration, and the lowest mean design label of configurations selected for that $\tau$). Results on or above the dashed diagonal line indicate that selected configurations are successful.
@@ E.2 Density Ratio Estimation
To estimate the density ratio function, $p_{X;\lambda}(\cdot)/p_{\text{lab}}(\cdot)$, for every configuration on the menu, $\lambda\in\Lambda$, we used multinomial logistic regression-based density ratio estimation (MDRE) (Srivastava et al.,, 2023). MDRE builds upon a formal connection between density ratio estimation (DRE) and classification (Bickel et al.,, 2009; Gutmann and Hyvärinen,, 2012): for the (correctly specified) binary classifier that minimizes the population cross-entropy risk in distinguishing between samples from two distributions, its logit for any input $x$ is equivalent to the density ratio for $x$, $p_{X;\lambda}(x)/p_{\text{lab}}(x)$ (Gutmann and Hyvärinen,, 2012). In practice, however, we can only hope to find a classifier that minimizes the empirical risk between finite samples from the two distributions. This distinction between the population and empirical risks hinders DRE far more than it does classification for classification’s sake: obtaining the density ratio requires getting the exact value of the population-optimal classifier’s logit, whereas optimal classification performance can be achieved by any classifier that learns the same decision boundary as the optimal classifier, even if its logits differ from the optimal classifier. Accordingly, we found that for many configurations on the menu, classifiers with very low training and validation losses often yielded poor approximations of the density ratio. This was particularly true when the design and labeled distributions were far apart, because the classification problem is too easy given finite samples: many different classifiers can minimize the empirical risk, none of which may happen to coincide with the population-optimal one whose logits are equivalent to density ratios. Telescoping density ratio estimation (Rhodes et al.,, 2020) tackles this problem by constructing intermediate distributions that interpolate between the numerator and denominator distributions, creating a sequence of “harder” binary classification problems for which there are fewer empirically optimal classifiers, and whose resulting estimated density ratios can be combined through a telescoping sum to approximate the original density ratio of interest. MDRE is similarly motivated but constructs a single multi-class classification problem between the intermediate distributions, justified by theoretical connections between the population-optimal classifier and the density ratio analogous to those described above (Srivastava et al.,, 2023). Concretely, let $h^{c}(x)$ denote the unnormalized log-probability according to a trained classifier that $x$ belongs to distribution $c\in\{1,\ldots,C\}$, where $C$ denotes the total number of distributions. MDRE uses $\exp(h^{i}(x)-h^{j}(x))$ to approximate the density ratio between distributions $i$ (numerator) and $j$ (denominator).